OpenStudy (anonymous):

Simplify: 4p+9+(-7p)+2=?

OpenStudy (anonymous):

4p-7p+9+2=?

OpenStudy (anonymous):

4p-7p=-3p and 9+2=11 so it's -3p+11

OpenStudy (anonymous):

okay, that one i had, i just wanted to make sure i had it correct. thank you!